# CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation

### v2

@@ Abstract
Large language models (LLMs) have revolutionized natural language processing (NLP), particularly through Retrieval-Augmented Generation (RAG), which enhances LLM capabilities by integrating external knowledge. However, traditional RAG systems face critical limitations, including disrupted contextual integrity due to text chunking, and over-reliance on semantic similarity for retrieval. To address these issues, we propose CausalRAG, a novel framework that incorporates causal graphs into the retrieval process. By constructing and tracing causal relationships, CausalRAG preserves contextual continuity and improves retrieval precision, leading to more accurate and interpretable responses. We evaluate CausalRAG against regular RAG and graph-based RAG approaches, demonstrating its superiority across several metrics. Our findings suggest that grounding retrieval in causal reasoning provides a promising approach to knowledge-intensive tasks.
@@ 1 Introduction
Large language models (LLMs) have transformed natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024). However, their reliance on static, pre-trained knowledge limits their ability to incorporate and reason over dynamically updated external information, particularly in knowledge-intensive domains. Retrieval-Augmented Generation (RAG) addresses this limitation by combining external retrieval with generative modeling to improve contextual understanding and response quality Lewis et al. (2021).
@@
Recent efforts to improve RAG focus on two fronts: 1) enhancing retrieval efficiency through adaptive and modular frameworks Gan et al. (2024); Ravuru et al. (2024); Zhang et al. (2024a); and 2) better structuring external knowledge, with graph-based RAGs emerging as a dominant approach Edge et al. (2024); Guo et al. (2024); Potts (2024). Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
@@
Through theoretical and empirical analysis, we rethink the limitations of current RAG systems by introducing a novel perspective grounded in context recall and precision metrics. Using this lens, we find that both regular and graph-based RAGs often fail to retrieve causally grounded content or accurately align it with the user query. We identify this fundamental issue as a primary reason why LLMs in RAG frameworks frequently produce seemingly relevant yet shallow responses that lack essential details.
@@
To address these gaps, we propose CausalRAG, a novel RAG framework that integrates causal graphs to guide retrieval. By identifying cause–effect relationships within external knowledge, CausalRAG preserves contextual coherence and improves reasoning fidelity. This leads to more accurate and causally grounded responses, while reducing hallucinations and enhancing answer faithfulness.
@@
We evaluate CausalRAG across datasets and varying context lengths, comparing it with regular and other competitive graph-based RAGs over multiple metrics. Results demonstrate that CausalRAG achieves superior performance across different contexts. Additionally, we conduct a case study and a parameter analysis to further examine our framework, analyzing and providing insights that contribute to ongoing research in RAG. The contributions of this work are threefold:
  • We identify the root limitations of current RAG systems through analytical and empirical studies, revealing why LLMs often produce superficial and under-grounded responses.
  • We propose CausalRAG, which improves retrieval and generation by grounding responses in causal structure.
  • Our work further mitigates hallucination issues and enhance interpretability, offering new insights into the design of RAG systems.
@@ 2.1 Retrieval-Augmented Generation
RAG enhances LLMs’ ability to handle knowledge-intensive tasks by integrating external knowledge retrieval Lewis et al. (2021). Existing research advances RAG in two main directions: 1) optimizing retrieval flow and interaction; and 2) structuring and utilizing external knowledge—particularly through knowledge graphs—for multi-stage reasoning.
Optimizing retrieval flow and interaction. The first stream focuses on refining retrieval dynamics to improve output quality. Several methods introduce pre-, mid-, and post-retrieval enhancements to reduce redundancy and computation Wang et al. (2024). Modular RAG architectures enable iterative retrieval–generation loops for more adaptive workflows. For instance, CAiRE-COVID Su et al. (2020) demonstrated multi-document summarization with iterative retrieval, and others extended this to multi-hop QA Feng et al. (2023). Recent work like METRAG Gan et al. (2024) leverages LLM supervision for utility-driven retrieval, while RAFT Zhang et al. (2024a) trains models to disregard distractor documents through chain-of-thought reasoning.
Structuring and utilizing external knowledge for multi-stage reasoning. The second stream focuses on representing and leveraging external knowledge—often via knowledge graphs—to support multi-level reasoning. GraphRAG Edge et al. (2024) represents documents as nodes and edges and summarizes subgraph communities for retrieval. LightRAG Guo et al. (2024) introduces dual-level retrieval and supports graph dynamic updates, while Lazy GraphRAG Potts (2024) further improves efficiency by delaying heavy computations until query time.
More recent multi-stage RAGs formalize hierarchical reasoning across multiple stages. PolyRAG Chen et al. (2025) uses a knowledge pyramid for coarse-to-fine retrieval. KG2RAG Zhu et al. (2025) guides RAG with KG-guided chunk expansion process to enhance factual grounding and control. GenTKGQA Gao et al. (2024) performs two-stage retrieval and generation over temporal subgraphs. HippoRAG2 Gutiérrez et al. (2025) integrates PageRank ranking with deeper passage reasoning for better answer quality. However, most of these approaches focus primarily on matching and ranking information within the knowledge graph, with limited attention to the causal intent behind the user’s query. A key challenge in RAG remains ensuring that retrieved information is not only relevant but also coherently aligned with the user’s underlying reasoning needs and the generation objective Gupta et al. (2024).
@@ 2.2 Causal Graphs and RAG
Causal graph integration has emerged as a promising way to improve retrieval and reasoning. Causal structures support more interpretable and reliable outputs by modeling underlying dependencies Ma (2024). Prior work in this area primarily focuses on causal discovery using RAG and LLMs. For instance, LLM-assisted BFS has been used to accelerate full causal graph discovery Jiralerspong et al. (2024), while Corr2Cause evaluates LLMs’ ability to infer causation from correlation Jin et al. (2024).
Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al. (2024), whereas the direct integration of causal graphs into RAG architectures remains largely unexplored. Our work aims to be a pioneer in this direction. A few existing studies have touched upon this concept but differ in scope. One integrates causality within the transformer architecture itself Chen et al. (2024), and another uses causal graphs only in pre-retrieval stages without deeper exploration Samarajeewa et al. (2024). GraphRAG Edge et al. (2024), while influential for its use of community detection and graph summarization, does not incorporate causality—nor do many recent multi-stage graph-based RAG frameworks.
In the following sections, we first analyze the limitations of regular and graph-based RAGs using a new recall–precision lens. We then introduce a causal graph structure to address these gaps.
@@ 3.1 Limitations of Regular RAG
Context $\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
@@
Context $\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
@@ 3.2 Rethinking Graph-based RAGs
Experimental study. Figure 1(b) reports experimental results for Regular RAG and both variants of GraphRAG. The graph-based approaches markedly improve context precision, reflecting the additional reranking and subgraph-summary steps in their retrieval pipelines. This gain comes with a modest reduction in context recall, indicating that some relevant passages are pruned during summarization. These scores were obtained with the Ragas evaluation framework Es et al. (2023); more comprehensive experiments and implementation details are shown in Section 5.
@@ 4.2 Discovering and Estimating Causal Paths
At query time, we first match the user query to graph nodes using embedding distance and select the top-$k$ closest nodes. The parameter $k$ controls the retrieval breadth: higher values return more context but increase computational cost. We then expand these nodes along graph edges by a step size $s$, which allows the framework to preserve causally and relationally connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall. The parameter $s$ determines the expansion depth and diversity of retrieved information.
Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph and generating the causal summary report (see LLM prompts in Appendix A). LLMs have demonstrated superiority in discerning and analyzing causal relationships Zhang et al. (2024b); Zhou et al. (2024) and this step ensures that CausalRAG prioritizes causally relevant information, improving precision. The resulting causal graph serves two purposes: (1) capturing long-range cause–effect dependencies often missed by semantic retrieval, particularly in longer texts; and (2) filtering out superficially related but causally irrelevant content, reducing hallucinations and improving faithfulness.
@@ 4.3 Retrieving Context Causally
After tracing and identifying key causal paths in the causal summary, we feed the result into the next stage and combine it with the user’s query (see LLM prompts in Appendix A). The causal summary prioritizes nodes that directly support the query, ensuring coherence and factual consistency while filtering out spurious information. This step reduces the risk of retrieving misleading yet superficially similar contexts. Finally, the combined prompt—including the causal summary—is used for generation, enabling CausalRAG to reason over major causal relationships rather than merely aggregating loosely related text spans.
@@ 5.1 Experimental Setup
Datasets. While many open-domain QA benchmark datasets exist, most are designed for explicit fact retrieval (e.g., “When was Google founded?” “1998”) and target classic NLP tasks. These datasets often fall short in evaluating discourse-level understanding, such as querying the underlying ideas, logic, or narrative within a document—tasks that better reflect real-world needs. Although some reading comprehension datasets exist Rajpurkar et al. (2018), their answers are typically short entities and designed for NLP models rather than RAG systems.
@@
Metrics & implementation details. We use the Ragas evaluation framework Es et al. (2023) to assess all models on three metrics: answer faithfulness, context precision, and context recall. The definitions of context precision and context recall are provided in the previous section. Answer faithfulness measures factual consistency on a scale from 0 to 100, with higher scores indicating closer alignment with reference documents. We use GPT-4o-mini as the base LLM for all frameworks. We set the parameters $k=s=3$ for CausalRAG, and use the same $k$ value for GraphRAG’s community-based retrieval, Regular RAG’s document retrieval, and HippoRAG2’s triple-based retrieval to ensure a fair comparison.
@@ 5.2 Performance Comparison
Causality delivers the most balanced and accurate retrieval. The model incorporating causal reasoning achieves the highest performance across all three metrics, with a context–precision score of $92.86$ and a faithfulness score of $78.00$, while maintaining competitive recall ($49.46$). By grounding retrieval in explicit causal paths, it filters out semantically similar but logically irrelevant content, thereby reducing superficial answers and improving the alignment and coherence between the retrieved context and the user query.
Graph-level structuring improves precision but introduces a recall–coverage trade-off. GraphRAG’s two operating modes illustrate this tension. Local search, which queries the raw knowledge graph, yields strong precision ($89.18$) but discards many supporting passages, dropping recall to $41.54$. Global search restores some of that missing evidence by summarizing community subgraphs, yet the broader scope simultaneously re-introduces off-topic material, pulling precision down to $66.67$ and faithfulness below $55.3$. These results corroborate Edge et al. (2024)’s observation that finer-grained community answers excel at factual grounding, whereas higher-level summaries aid breadth but dilute specificity.
Entity-centric multi-stage KGs narrow the gap but still overlook discourse context. HippoRAG2 augments Personalized PageRank with LLM-filtered triples and passage nodes, boosting precision to $73.72$ and faithfulness to $67.36$ without sacrificing recall relative to Regular RAG. Nevertheless, its reliance on entity extraction leaves many context-rich sentences unlinked to the query, a limitation they noted as concept–context trade-off and need for deeper contextualizationGutiérrez et al. (2025). By contrast, CausalRAG’s path-based expansion integrates both entities and their causal explanations, yielding a retrieval set that is simultaneously broader in coverage and more tightly aligned with the question, thereby establishing a new possibility for knowledge-intensive RAG tasks.

### v3

@@ Abstract
Large language models (LLMs) have revolutionized natural language processing (NLP), particularly through Retrieval-Augmented Generation (RAG), which enhances LLM capabilities by integrating external knowledge. However, traditional RAG systems face critical limitations, including disrupted contextual integrity due to text chunking, and over-reliance on semantic similarity for retrieval. To address these issues, we propose CausalRAG, a novel framework that incorporates causal graphs into the retrieval process. By constructing and tracing causal relationships, CausalRAG preserves contextual continuity and improves retrieval precision, leading to more accurate and interpretable responses. We evaluate CausalRAG against regular RAG and graph-based RAG approaches, demonstrating its superiority across several metrics. Our findings suggest that grounding retrieval in causal reasoning provides a promising approach to knowledge-intensive tasks [^1] Code is publicly available at https://github.com/Pwnb/CausalRAG..
@@ 1 Introduction
The rapid advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024). However, their reliance on pre-trained knowledge limits their ability to integrate and reason over dynamically updated external information, particularly in knowledge intensive domains such as academic research. Retrieval-Augmented Generation (RAG) has emerged as a promising framework to address this limitation Lewis et al. (2021), combining retrieval mechanisms with generative capabilities to enhance contextual understanding and response quality.
@@
Recent research has focused on improving RAG along two primary directions: 1) enhancing retrieval efficiency and integration mechanisms by designing more adaptive and dynamic retrieval frameworks Gan et al. (2024); Ravuru et al. (2024); Zhang et al. (2024a); 2) improving the representation of external knowledge to facilitate retrieval and reasoning, with graph-based RAGs being a dominant approach Edge et al. (2024); Guo et al. (2024); Potts (2024). Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
@@
Through a combination of theoretical analysis and empirical evaluation, we rethink the limitations of current RAG systems by introducing a novel perspective based on context recall and precision metrics. Our findings reveal that both regular and graph-based RAGs struggle not only to retrieve truly grounded context but also to accurately discern the relationship between retrieved content and the user query. We identify this fundamental issue as one primary reason why LLMs in RAG frameworks often generate seemingly relevant yet shallow responses that lack essential details.
@@
To address these gaps, we introduce CausalRAG, a novel RAG framework that integrates causal graphs to enhance retrieval accuracy and reasoning performance. Unlike regular and graph-based RAGs, CausalRAG explicitly identifies causal relationships within external knowledge, preserving contextual coherence while capturing underlying cause–effect dependencies. By ensuring that retrieved documents are both relevant and causally grounded, CausalRAG enables the generation of more contextually rich and causally detailed responses. This approach not only improves retrieval effectiveness but also mitigates hallucinations and enhances answer faithfulness.
@@
We evaluate CausalRAG on datasets from diverse domains and across varying context lengths, comparing its performance with regular RAG and other graph-based RAG frameworks. Our experiments assess performance across three key metrics: answer faithfulness, context recall, and context precision. Results demonstrate that CausalRAG achieves superior performance across different contexts. Additionally, we conduct a case study and a parameter analysis to further examine our framework, analyzing and providing insights that contribute to ongoing research in RAG. The contributions of this work are threefold:
  • We systematically identify the inherent limitations of RAG’s retrieval process through analytical and experimental study. We uncovered the major reason why LLMs in RAG tend to generate superficial, generalized answers that lack the grounded details expected by users.
  • We propose CausalRAG, a framework that enhances both retrieval and generation quality by incorporating causality into the RAG, effectively addressing these limitations.
  • Our work further mitigates hallucination issues and significantly improves the interpretability of AI systems. We summarize key findings and insights in both retrieval and generation process, contributing to RAG research.
@@ 2.1 Retrieval-Augmented Generation
RAG enhances LLMs’ ability to handle knowledge-intensive tasks by integrating external knowledge retrieval Lewis et al. (2021). Existing research primarily advances RAG along two key dimensions: 1) improving retrieval efficiency and integration mechanisms; 2) enhancing the representation and multi-stage utilization of external knowledge—particularly through knowledge graphs—to support deeper reasoning and generation.
Optimizing Retrieval Flow and Interaction. The first stream focuses on improving the interaction flow within the RAG system to enhance output quality. Approaches have introduced pre-retrieval, retrieval, and post-retrieval refinements to mitigate redundancy and computational overhead Wang et al. (2024). Modular RAG architectures further advance this by enabling iterative retrieval-generation cycles, allowing dynamic interactions between retrieval and content creation. For example, CAiRE-COVID Su et al. (2020) demonstrated the effectiveness of iterative retrieval in multi-document summarization, while some work Feng et al. (2023) extended this approach to multi-hop question answering. Recent innovations include METRAG Gan et al. (2024), which integrates LLM supervision to generate utility-driven retrieval processes, and RAFT Zhang et al. (2024a), which trains models to disregard distractor documents while improving citation accuracy through chain-of-thought reasoning.
Structuring and Utilizing External Knowledge for Multi-stage Reasoning. The second stream focuses on representing and processing external knowledge—often in the form of knowledge graphs—to improve retrieval depth and reasoning accuracy. GraphRAG Edge et al. (2024) treats documents as interconnected nodes to capture thematic and causal links, while LightRAG Guo et al. (2024) adopts a lightweight, dual-level retrieval approach for dynamic updates. Lazy GraphRAG Potts (2024) further improves efficiency by delaying heavy computations until query time.
More recently, several multi-stage RAG frameworks have emerged that explicitly structure the retrieval and generation process across multiple levels. PolyRAG Chen et al. (2025) builds a hierarchical knowledge pyramid, allowing coarse-to-fine retrieval over different abstraction layers. KG2RAG Zhu et al. (2025) guides RAG with KG-guided chunk expansion process to enhance factual grounding and control. GenTKGQA Gao et al. (2024) employs a two-stage pipeline to first retrieve temporal subgraphs and then generate based on virtual knowledge indicators. Among them, HippoRAG2 Gutiérrez et al. (2025) focuses on combining PageRank algorithm with deeper passage integration, improving both knowledge retention and answer quality. However, most of these approaches focus primarily on matching and ranking information within the knowledge graph, with limited attention to the causal intent behind the user’s query. A key challenge in RAG remains ensuring that retrieved information is not only relevant but also coherently aligned with the user’s underlying reasoning needs and the generation objective Gupta et al. (2024).
@@ 2.2 Causal Graphs and RAG
The combination of causal graphs and RAG has emerged as a promising approach to enhancing knowledge retrieval and reasoning. As causality provides a structural understanding of dependencies within data, it enables more interpretable and reliable AI outputs Ma (2024). Existing research in this domain primarily advances causal discovery with RAG and LLMs. For instance, some work proposed an LLM-assisted breadth-first search (BFS) method for full causal graph discovery, significantly reducing time complexity Jiralerspong et al. (2024). Additionally, some further introduced a correlation-to-causation inference (Corr2Cause) task to evaluate LLMs’ ability to infer causation from correlation, revealing their limitations in generalization across different datasets (Jin et al., 2024).
Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al. (2024), whereas the direct integration of causal graphs into RAG architectures remains largely unexplored. Our work aims to be a pioneer in this direction. A few existing studies have touched upon this concept but differ in scope. One approach integrates causal graphs within the LLM architecture itself, structuring the transformer’s internal token processing using causality rather than enhancing RAG retrieval Chen et al. (2024). Another employs causal graphs in RAG systems but focuses on the pre-retrieval stage and largely reduces the core process into a single embedding model without deeper exploration Samarajeewa et al. (2024). GraphRAG Edge et al. (2024), while influential for its use of community detection and graph summarization, does not incorporate causality—nor do many recent multi-stage graph-based RAG frameworks.
In the following sections, we first analyze the nature of regular RAG and grah-based RAGs from a novel perspective, identifying their inherent limitations. We then introduce a causal graph structure to address these gaps.
@@ 3.1 Limitations of Regular RAG
$\text{Context Recall}=\frac{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}{|R|}$ (1)
@@
$\text{Context Precision}=\frac{\sum_{i=1}^{N}\mathbb{I_{Q}}(C_{i}\in R)}{\sum_{i=1}^{N}\mathbb{I}(C_{i}\in R)}$ (2)
@@ 3.2 Rethinking Graph-based RAGs
Experimental Study. Figure 1(b) reports experimental results for Regular RAG and both variants of GraphRAG. The graph‑based approaches markedly improve context precision, reflecting the additional reranking and subgraph‑summary steps in their retrieval pipelines. This gain comes with a modest reduction in context recall, indicating that some relevant passages are pruned during summarization. These scores were obtained with the Ragas evaluation framework Es et al. (2023); more comprehensive experiments and implementation details are shown in Section 5.
@@ 4.2 Discovering and Estimating Causal Paths
This is the first step during query time. We begin by matching the user query to nodes in the graph based on their embedding distance. The $k$ nodes with the smallest distances are selected, representing the most relevant information directly related to the query. Notably, $k$ is a tunable parameter—higher values retrieve more relevant information at the cost of increased computational complexity.
After selecting the initial $k$ nodes, we expand the search along the base graph’s edges by a step size of $s$, thereby broadening the retrieved context. This step is crucial as it preserves causal and relational connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall. The parameter $s$ controls the depth of expansion, where higher values lead to more diverse information retrieval.
Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph and generating the causal summary report (see LLM prompts in Appendix A). LLMs have demonstrated superiority in discerning and analyzing causal relationships Zhang et al. (2024b); Zhou et al. (2024) and this step ensures that CausalRAG prioritizes causally relevant information, improving precision.
Furthermore, the derived causal graph serves two key purposes: 1) It preserves causally relevant information that traditional retrieval methods struggle to capture. More importantly, by adjusting the parameter $s$, this approach can capture long-range causal relationships within the text, particularly when the text is lengthy; 2) It filters out semantically related but causally irrelevant information. Without this filtering, responses may contain unnecessary or even hallucinated content, compromising answer faithfulness.
@@ 4.3 Retrieving Context Causally
After constructing the causal graph, we summarize the retrieved information and generate a causal summary (see LLM prompts in Appendix A). Notably, the input at this stage consists of information that is not only highly relevant but also causally grounded in the user’s query, ensuring greater validity. This approach contrasts with traditional retrieval methods, which often rely purely on semantic similarity and may retrieve contextually related yet causally irrelevant information.
The causal summary is derived by tracing key causal paths within the graph, prioritizing nodes and relationships that contribute directly to answering the query. This ensures that the retrieved information maintains logical coherence and factual consistency while filtering out spurious or weakly related context. Additionally, by leveraging causal dependencies, our method reduces the risk of retrieving semantically similar but misleading evidence.
Once the causal summary is generated, it is combined with the user query to construct a refined prompt for CausalRAG. This structured final input allows RAG to focus on reasoning through causal relationships rather than merely aggregating loosely related text spans.
@@ 5.1 Experimental Setup
Datasets. While many open-domain QA benchmark datasets exist, most are designed for explicit fact retrieval (e.g., “When was Google founded?” “1998.”) and target classic NLP tasks. These datasets often fall short in evaluating discourse-level understanding, such as querying the underlying ideas, logic, or narrative within a document—tasks that better reflect real-world needs. Although some reading comprehension datasets exist Rajpurkar et al. (2018), their answers are typically short entities and designed for NLP models rather than RAG systems.
@@
Metrics and Implementation Details. We use the Ragas evaluation framework Es et al. (2023) to assess all models on three metrics: answer faithfulness, context precision, and context recall. The definitions of context precision and context recall are provided in the previous section. Answer faithfulness measures factual consistency on a scale from 0 to 100, with higher scores indicating closer alignment with reference documents. We use GPT-4o-mini as the base LLM for all frameworks. We set the parameters $k=s=3$ for CausalRAG, and use the same $k$ value for GraphRAG’s community-based retrieval, Regular RAG’s document retrieval, and HippoRAG2’s triple-based retrieval to ensure a fair comparison.
@@ 5.2 Performance Comparison
Causality delivers the most balanced and accurate retrieval. The experimental results, as depicted in Figure 3, reveal distinct strengths and weaknesses among the evaluated RAG frameworks, ultimately highlighting the advantages of integrating causal reasoning. CausalRAG consistently achieves the most balanced and superior performance, leading in answer faithfulness ($78.00$), context precision ($92.86$), and maintaining competitive context recall ($49.46$). This suggests that by explicitly tracing and retrieving along causal paths, CausalRAG effectively mitigates the common RAG pitfalls of relying purely on semantic similarity and grappling with contextual integrity.
Graph‑level structuring improves precision but introduces a recall–coverage trade‑off. Graph-based RAG approaches, while an improvement over regular RAG in certain aspects, demonstrate a clear trade-off. GraphRAG-Local excels in context precision ($89.18$) but suffers from a low context recall ($41.54$). This indicates that while its direct querying of the raw knowledge graph yields highly relevant local information, it often discards broader supporting passages. The global variant, GraphRAG-Global, attempts to address this by summarizing community subgraphs, thereby improving recall ($47.22$). However, this comes at a significant cost to precision ($66.67$) and faithfulness ($55.27$). These results corroborate Edge et al. (2024)’s observation that finer‑grained community answers excel at factual grounding, whereas higher‑level summaries aid breadth but dilute specificity. In contrast, CausalRAG manages to maintain high precision while achieving better recall through its targeted causal path selection.
Entity‑centric multi‑stage KGs narrow the gap but still overlook discourse context. HippoRAG2 presents a strong performance, particularly in answer faithfulness ($67.36$) and context precision ($73.72$) when compared to Regular RAG and GraphRAG-Global, with a recall of $47.22$. Its methodology, which augments Personalized PageRank with LLM-filtered triples and passage nodes, clearly enhances retrieval quality. Nevertheless, its reliance on entity extraction leaves many context‑rich sentences unlinked to the query, a limitation they noted as concept–context trade‑off and need for deeper contextualizationGutiérrez et al. (2025). Then we found CausalRAG surpasses it across all metrics. By comparison, CausalRAG’s causal path expansion allows for the integration of both entities and their underlying causal explanations. This approach yields a retrieval set that is not only broad in coverage but also more tightly aligned with the user’s query, leading to more accurate and faithful responses by ensuring the retrieved context is causally pertinent rather than just semantically adjacent.